# Controlling Impact Noise Propagation With Ceramic Metatiles

### v1

@@ II.3 Numerical results
Secondly, for the variation of the silicon joint width ($c$), we fixed the width of the subtiles at $b=5$ cm and the Young’s modulus of silicon at $1$ MPa, as depicted in Fig. 2(b). Then, we varied $c$ from $0.1$ cm to $0.5$ cm to limit the size of silicon joints, ensuring that the ceramic metatile retains its resemblance to a standard one, with only small additional silicon joints. It is observed that for $c=0.1$ cm, the bandgap starts at $831$ Hz. As the width of the silicon joints increases, this bandgap shifts toward lower frequencies, reaching $328$ Hz for $c=0.5$ cm. This parametric study provides insights into how the silicon joints influence the frequency range of the bandgaps and how they can be controlled by introducing silicon joints into the ceramic metatile. However, the dimensions of the silicon joints also depend on the target frequency range for the intended application.
@@
In the final parametric study, we fixed the dimensions of the ceramic metatile with a subtile width of $5$ cm and a silicon joint width of $0.5$ cm, as depicted in Fig. 2(c). We varied the Young’s modulus of silicon to better understand how silicon mechanical properties affect noise cancellation. By keeping the density and Poisson’s ratio of the silicon constant, we varied the Young’s modulus from $1$ MPa to $10$ MPa. It is observed that as the Young’s modulus decreases, the bandgaps shift to lower frequencies. For example, for $E=1$ MPa, the bandgaps start at $328$ Hz. As the silicon becomes stiffer, the bandgaps shift to higher frequencies. For $E=10$ MPa, the bandgaps start at $1037$ Hz.
@@
This analysis of the material and geometric parameters provides us with a better understanding of the bandgaps and their appearance in the ceramic metatile. After this detailed analysis, we chose $c=0.5$ cm, $b=5$ cm, and a silicon Young’s modulus of $1$ MPa, which we consider for the rest of the paper. Using these values, we plot the dispersion curves in the first irreducible Brillouin zone along the $[100]$ direction. The results of this study are presented in Fig. 3. This figure shows the bandgaps, represented by the gray-colored region, and the propagating modes, which are plotted in blue. Additionally, we display the eigenmodes of the ceramic metatile around the bandgaps, shown as screenshots of the total modulus displacement distribution for a ceramic metatile composed of $3$ unit cells.
Figure 4: Numerical transmission curves for transverse waves. (a) Normal to the silicon joints along the $[100]$ direction. (b) Diagonal to the silicon joints along the $[110]$ direction. (c) Parallel to the silicon joints along the $[010]$ direction.
After calculating the phononic dispersion of the ceramic metatile, modeled as an infinite structure, we now proceed to characterize its dynamic behavior by analyzing the transmission of elastic waves through a single ceramic tile. This tile consists of $5$ silicon joints and $6$ ceramic subtiles, as illustrated in Fig. 1(c). The transmission spectrum is calculated along the three principal directions: $[100]$, $[110]$, and $[010]$, to assess the ceramic metatile’s response to wave propagation in different orientations.
@@
We begin by applying a transverse wave and calculating its transmission, as illustrated in Fig. 4. Fig. 4(a) presents the computed transverse transmission along the $[100]$ direction, where the red curve corresponds to the ceramic metatile, the blue curve represents the standard ceramic tile and the black curve illustrates the transmission through one silicon joint. In this direction, the transmissions of the standard ceramic tile and the $1$ silicon joint remain high across the entire frequency range studied ($20-5000$ Hz). In contrast, the ceramic metatile exhibits a significant reduction in transmission within the $500-1900$ Hz range, whereas in the lower range ($20-500$ Hz), the transmission remains nearly total. Beyond $1900$ Hz, the bandgaps are affected by diffraction limits, yet they retain their general shape and continue to attenuate vibrations effectively.
@@
In the $[110]$ direction (Fig. 4(b)), the bandgap extends from $500$ Hz to $1400$ Hz, after which the transmission increases until $2000$ Hz. Beyond this point, additional bandgaps appear, influenced by the diffraction limit. Meanwhile, the standard ceramic tile maintains nearly maximal transmission across the entire frequency spectrum. When analyzing transmission in the $[010]$ direction (Fig. 4(c)), where wave propagation is parallel to the silicon joints, we observe that the ceramic metatile’s transmission closely resembles that of the standard ceramic tile. This indicates that the presence of silicon joints in the propagation direction of transverse waves attenuates vibrations similarly to the inclined $45\textdegree$ joints. These findings suggest that our design effectively reduces transverse waves propagation in directions where the silicon joints are not aligned parallel to the wave propagation.
@@
To further validate the dispersion curves in the first irreducible Brillouin zone for all existing modes, we now compute the transmission for a longitudinally polarized wave. Using the same model, we apply a longitudinal wave as the input displacement and measure the output displacement. The results of this study are illustrated in Fig. 5, where the transmission of the longitudinal wave is analyzed in three different directions. Fig. 5(a) shows the transmission along the $[100]$ direction. The standard ceramic tile exhibits excellent transmission of longitudinal vibrations, whereas the ceramic metatile displays significantly reduced transmission between $400$ Hz and $1950$ Hz. Beyond this range, bandgaps emerge, coupled with diffraction limits. Furthermore, we analyze the behavior of the inclined joints in relation to longitudinal wave propagation along the $[110]$ direction. Our design demonstrates a broad bandgap extending from $400$ Hz to $1950$ Hz, followed by the presence of bandgaps at higher frequencies, again influenced by diffraction effects, ad shown in Fig. 5(b). Finally, in the $[010]$ direction, the transmission behavior closely resembles that observed for transverse waves. In this case, both the standard ceramic tile and the ceramic metatile exhibit high transmission for both longitudinal and transverse waves, as depicted in Fig. 5(c).
@@
We numerically demonstrate that our design, which incorporates silicon joints into ceramic metatiles, exhibits broadband bandgaps at low frequencies. Moreover, these bandgaps can be controlled by adjusting the geometrical and material parameters. The dispersion calculations presented in this section confirm that our chosen parameters create four distinct bandgaps in the low-frequency range of $20-2000$ Hz for an infinitely periodic structure. To further validate the presence of these bandgaps in a finite structure, we calculated the transmission of both transverse and longitudinal waves through a single ceramic metatile. Additionally, we numerically compared its response to that of a standard ceramic tile. In the next section, we present the experimental validation of these bandgaps.
III Experimental validation
III.1 Laboratory transmission measurements
@@
To validate numerical results, a ceramic metatile was fabricated by assembling ceramic plates with silicon joints, as shown in Fig. 6(c). The fabricated metatile has geometric parameters of width of $32.5cm$ and $30cm$, with a height of $1cm$, which is composed of small ceramic plates of $5cm$ as width and separated by silicon joint with a width of $0.5cm$.
@@
Fig. 6(a) illustrates the experimental setup used to measure the transmission spectrum of the ceramic metatile. A signal generator was employed to produce a sweep signal, which was fed into a power amplifier to enhance the signal amplitude. The amplified signal was then transmitted to a PZT patch, which generated the input displacement. A transducer measured this displacement, while another transducer at the output recorded the transmitted signal. The output transducer was connected to an analyzer, which interfaced with a laptop for data acquisition. This setup enabled the measurement and recording of both transmitted and incident displacements, allowing for the calculation of transmission curves.
@@
Figure 7: Experimental transmission curves for transverse waves. (a) Normal to the silicon joints along the $[100]$ direction. (b) Diagonal to the silicon joints along the $[110]$ direction. (c) Parallel to the silicon joints along the $[010]$ direction.
III.2 Experimental results
@@
For the experimental transmission measurements, we analyze the propagation of transverse waves along three directions: $[100]$, $[010]$, and $[110]$. Then, we compare the results obtained from a standard ceramic tile with those of our designed ceramic metatile. Fig. 7 presents the experimental results, where the blue curves correspond to the standard ceramic tile, and the red curves represent the ceramic meta-tile’s response in all directions. In the $[100]$ direction (Fig. 7(a)), the transmission remains high between $20$ Hz and $700$ Hz. Beyond this range, the transmission of the metatile decreases by approximately $50$ dB compared to the standard tile from $700$ Hz to $2500$ Hz. For the $[110]$ direction (Fig. 7(b)), a bandgap appears between $700$ Hz and $1400$ Hz. The transmission then increases until $1490$ Hz, after which the metatile exhibits a transmission reduction of around $50$ dB relative to the standard ceramic tile. finally, for the $[010]$ direction (Fig. 7(a)), both tiles exhibit high transmission of transverse waves.
IV Discussion
@@
After numerically and experimentally demonstrating the existence of stop bandgaps in the ceramic metatile, this section focuses on studying impact noise reduction in both the standard ceramic tile and the metatile using a real impact noise source. The setup, illustrated in Fig. 8(a) and (c), involves a shaker connected to a hammer-like object to simulate the impact noise generated by footsteps on tiled floors. We excite the shaker by a theoretical signal, which is a Gaussian-like signal with an interval of $100$ ms and we employ the same laboratory setup as depicted in Fig. 6.
@@
A transducer is placed to measure the impact noise source, while another transducer is placed on the opposite side of the tiles in order to measure their transmission noise amplitude. The results of this study are shown in Fig. 8 (b) and (d), which show the time-dependent acceleration amplitude for both the standard ceramic tile and the metatile. The impact noise source is represented by the blue curve, while the transmitted noise is shown in red. Standard ceramic tile exhibits high transmitted noise, characterized by a large amplitude over time. In contrast, the metatile shows a significantly lower amplitude, indicating its effectiveness in attenuating noise propagation.
@@
To further analyze this behavior, the Fast Fourier Transform (FFT) of both responses is calculated and presented in Fig. 8(e) and (f). The standard ceramic tile allows noise transmission across the frequency spectrum from $20$ Hz to $2000$ Hz. However, the metatile demonstrates a significant reduction in noise transmission at low frequencies, and beyond $500$ Hz, no impact noise transmission is observed. These results highlight the strong noise control and attenuation capabilities of the metatile, making it a promising solution for tiled flooring applications.
V Conclusion
In conclusion, we have proposed a novel design of metatiles that combines ceramic tiles with silicon joints. This design highlights the essential properties of ceramic tiles in noise reduction as tiled floors. Broadband bandgaps at low frequency ranges are obtained for both transverse and longitudinal waves. For transverse waves, the ceramic metatile exhibits a significant reduction in transmission within the $500-1900$ Hz range along the $[100]$ direction and within the $500-1400$ Hz range along the $[110]$ direction, respectively. For longitudinal waves, the ceramic metatile displays a broad bandgap extending from 400 Hz to 1950 Hz along the $[100]$ and $[110]$ directions. Besides, the bandgaps shift towards lower frequencies as the width of subtiles and silicon joints increases, or as the Young’s modulus of silicon decreases, which indicates that they can be tailored by geometry parameters and material properties. Furthermore, we fabricate metatile samples and experimentally validate their strong noise control and attenuation capabilities, showcasing their potential for practical use in architectural acoustics for tiled floors.

### v2

@@ 2.3 Numerical results
Secondly, for the variation of the silicon joint width $c$, we fixed the width of the subtiles at $b=5$ cm and the Young’s modulus of silicon at $1$ MPa, as depicted in Figure 2(b). Then, we varied $c$ from $0.1$ cm to $0.5$ cm to limit the size of silicon joints, ensuring that the ceramic metatile retains its resemblance to a standard one, with only small additional silicon joints. It is observed that for $c=0.1$ cm, the bandgap starts at $831$ Hz. As the width of the silicon joints increases, this bandgap shifts toward lower frequencies, reaching $328$ Hz for $c=0.5$ cm. This parametric study provides insights into how the silicon joints influence the frequency range of the bandgaps and how they can be controlled by introducing silicon joints into the ceramic metatile. However, the dimensions of the silicon joints also depend on the target frequency range for the intended application.
@@
In the final parametric study, we fixed the dimensions of the ceramic metatile with a subtile width of $5$ cm and a silicon joint width of $0.5$ cm, as depicted in Figure 2(c). We varied the Young’s modulus of silicon to better understand how silicon mechanical properties affect noise cancellation. By keeping the density and Poisson’s ratio of the silicon constant, we varied the Young’s modulus from $1$ MPa to $10$ MPa. It is observed that as the Young’s modulus decreases, the bandgaps shift to lower frequencies. For example, for $E=1$ MPa, the bandgaps start at $328$ Hz. As the silicon becomes stiffer, the bandgaps shift to higher frequencies. For $E=10$ MPa, the bandgaps start at $1037$ Hz.
@@
This analysis of the material and geometric parameters provides us with a better understanding of the bandgaps and their appearance in the ceramic metatile. After this detailed analysis, we chose $c=0.5$ cm, $b=5$ cm, and a silicon Young’s modulus of $1$ MPa, which we consider for the rest of the paper. Using these values, we plot the dispersion curves in the first irreducible Brillouin zone along the $[100]$ crystallographic direction. The results of this study are presented in Figure 3. This figure shows the bandgaps, represented by the gray-colored region, and the propagating modes for the flexural waves, which are plotted in blue. Additionally, we display the eigenmodes of the ceramic metatile around the bandgaps, shown as screenshots of the total modulus displacement distribution for a ceramic metatile composed of $3$ unit cells. One can observe that the formation of the bandgaps affects all wave modes, meaning that within the bandgaps, no mode, whether flexural or longitudinal, can propagate. The first bandgap, starting from $F_{1}=332$ Hz to $F_{2}=860$ Hz, is due to Bragg scattering, which is related to the condition $\lambda=2\times b$, as depicted in the eigenmodes of the first-order mode vibration. Similarly, the second bandgap, from $F_{3}=882$ Hz to $F_{4}=1719$ Hz, corresponds to the eigenmode of the second-order vibration. In general, the bandgaps are formed by both Bragg scattering and the resonance of the mass-spring system created by the silicon joints and subtiles, which causes energy localization at the resonance frequencies of the system.
Figure 4: Numerical transmission curves for flexural waves. (a) Normal to the silicon joints along the $[100]$ direction. (b) Diagonal to the silicon joints along the $[110]$ direction. (c) Parallel to the silicon joints along the $[010]$ direction.
After calculating the phononic dispersion of the ceramic metatile, modeled as an infinite structure, we now proceed to characterize its dynamic behavior by analyzing the transmission of elastic waves through a single ceramic tile. This tile consists of $5$ silicon joints and $6$ ceramic subtiles, as illustrated in Figure 1(c). The transmission spectrum is calculated along the three principal directions: $[100]$, $[110]$, and $[010]$, to assess the ceramic metatile’s response to wave propagation in different orientations.
@@
We begin by applying a flexural wave and calculating its transmission, as illustrated in Figure 4. Figure 4(a) presents the computed flexural transmission along the $[100]$ direction, where the red curve corresponds to the ceramic metatile, the blue curve represents the standard ceramic tile and the black curve illustrates the transmission through one silicon joint. In this direction, the transmissions of the standard ceramic tile and the $1$ silicon joint remain high across the entire frequency range studied ($20-5000$ Hz). In contrast, the ceramic metatile exhibits a significant reduction in transmission within the $500-1900$ Hz range, whereas in the lower range ($20-500$ Hz), the transmission remains nearly total. Beyond $1900$ Hz, the bandgaps are affected by diffraction limits, yet they retain their general shape and continue to attenuate vibrations effectively.
@@
In the $[110]$ direction (Figure 4(b)), the bandgap extends from $500$ Hz to $1400$ Hz, after which the transmission increases until $2000$ Hz. Beyond this point, additional bandgaps appear, influenced by the diffraction limit. Meanwhile, the standard ceramic tile maintains nearly maximal transmission across the entire frequency spectrum. When analyzing transmission in the $[010]$ direction (Figure 4(c)), where wave propagation is parallel to the silicon joints, we observe that the ceramic metatile’s transmission closely resembles that of the standard ceramic tile. This indicates that the presence of silicon joints in the propagation direction of flexural waves attenuates vibrations similarly to the inclined $45\textdegree$ joints. These findings suggest that our design effectively reduces flexural waves propagation in directions where the silicon joints are not aligned parallel to the wave propagation.
@@
To further validate the dispersion curves in the first irreducible Brillouin zone for all existing modes, we now compute the transmission for a longitudinally polarized wave. Using the same model, we apply a longitudinal wave as the input displacement and measure the output displacement. The results of this study are illustrated in Figure 5, where the transmission of the longitudinal wave is analyzed in three different directions. Figure 5(a) shows the transmission along the $[100]$ direction. The standard ceramic tile exhibits excellent transmission of longitudinal vibrations, whereas the ceramic metatile displays significantly reduced transmission between $400$ Hz and $1950$ Hz. Beyond this range, bandgaps emerge, coupled with diffraction limits. Furthermore, we analyze the behavior of the inclined joints in relation to longitudinal wave propagation along the $[110]$ direction. Our design demonstrates a broad bandgap extending from $400$ Hz to $1950$ Hz, followed by the presence of bandgaps at higher frequencies, again influenced by diffraction effects, ad shown in Figure 5(b). Finally, in the $[010]$ direction, the transmission behavior closely resembles that observed for flexural waves. In this case, both the standard ceramic tile and the ceramic metatile exhibit high transmission for both longitudinal and flexural waves, as depicted in Figure 5(c).
@@
We numerically demonstrate that our design, which incorporates silicon joints into ceramic metatiles, exhibits broadband bandgaps at low frequencies. Moreover, these bandgaps can be controlled by adjusting the geometrical and material parameters. The dispersion calculations presented in this section confirm that our chosen parameters create four distinct bandgaps in the low-frequency range of $20-2000$ Hz for an infinitely periodic structure. To further validate the presence of these bandgaps in a finite structure, we calculated the transmission of both flexural and longitudinal waves through a single ceramic metatile. Additionally, we numerically compared its response to that of a standard ceramic tile. In the next section, we present the experimental validation of these bandgaps.
3 Experimental validation
This section presents the experimental setup for measuring the transmission of flexural waves through the ceramic metatiles, along with the results obtained using the Fast Fourier Transform (FFT).
3.1 Laboratory transmission measurements
@@
To validate numerical results, a ceramic metatile was fabricated by assembling ceramic plates with silicon joints, as shown in Figure 6(c). The fabricated metatile has geometric parameters of width of $32.5cm$ and $30cm$, with a height of $1cm$, which is composed of small ceramic plates of $5cm$ as width and separated by silicon joint with a width of $0.5cm$.
@@
Figure 6(a) illustrates the experimental setup used to measure the transmission spectrum of the ceramic metatile. A signal generator was employed to produce a sweep signal, which was fed into a power amplifier to enhance the signal amplitude. The amplified signal was then transmitted to a PZT patch, which generated the input displacement. A transducer measured this displacement, while another transducer at the output recorded the transmitted signal. The output transducer was connected to an analyzer, which interfaced with a laptop for data acquisition. This setup enabled the measurement and recording of both transmitted and incident displacements, allowing for the calculation of transmission curves.
@@
Figure 7: Experimental transmission curves for flexural waves. (a) Normal to the silicon joints along the $[100]$ direction. (b) Diagonal to the silicon joints along the $[110]$ direction. (c) Parallel to the silicon joints along the $[010]$ direction.
3.2 Experimental results
@@
For the experimental transmission measurements, we analyze the propagation of flexural waves along three directions: $[100]$, $[010]$, and $[110]$. Then, we compare the results obtained from a standard ceramic tile with those of our designed ceramic metatile. Figure 7 presents the experimental results, where the blue curves correspond to the standard ceramic tile, and the red curves represent the ceramic meta-tile’s response in all directions. In the $[100]$ direction (Figure 7(a)), the transmission remains high between $20$ Hz and $700$ Hz. Beyond this range, the transmission of the metatile decreases by approximately $50$ dB compared to the standard tile from $700$ Hz to $2500$ Hz. For the $[110]$ direction (Figure 7(b)), a bandgap appears between $700$ Hz and $1400$ Hz. The transmission then increases until $1490$ Hz, after which the metatile exhibits a transmission reduction of around $50$ dB relative to the standard ceramic tile. finally, for the $[010]$ direction (Figure 7(c)), both tiles exhibit high transmission of flexural waves. Furthermore, the experimental results show good agreement with the numerical predictions in terms of both frequency ranges and attenuation amplitude. This is demonstrated by the calculated absolute error between the two results, given by the equation: $Error=\frac{T_{exp}-T_{num}}{T_{exp}}$ in which $T_{exp}$ the transmission at normal incidence of the flexural wave, and $T_{num}$ is its counterpart computed by FEM (see Appendix A. In addition, we perform numerical simulations to investigate the effect of damping by introducing an imaginary component to the Young’s modulus of the silicon joint, incorporating the loss factor $\eta$ as follows: $E_{\text{silicon}}=E(1+i\eta)$. The transmission along the [100] direction is computed for three different damping scenarios in order to compare the numerical results with the experimental data (see Appendix B). We observe that increasing the damping in the silicon joint leads to broader band gaps and higher attenuation of the flexural waves.
To summarize, both the error analysis between experimental and numerical results, and the study of the silicon joint’s loss factor, confirm that the observed band gaps are associated with Bragg scattering and local resonances, and that the experimental results are in good agreement with the numerical predictions.
4 Discussion
After numerically and experimentally demonstrating the existence of stop bandgaps in the ceramic metatile, this section focuses on studying impact noise reduction in both the standard ceramic tile and the metatile using a real impact noise source. The setup, illustrated in Figure 8(a) and (c), involves a shaker connected to a hammer-like object to simulate the impact noise generated by footsteps on tiled floors.
@@ 4 Discussion
We excite the shaker by a theoretical signal, which is a Gaussian-like signal with an interval of $100$ ms and we employ the same laboratory setup as depicted in Figure 6.
@@
A transducer is placed to measure the impact noise source, while another transducer is placed on the opposite side of the tiles in order to measure their transmission noise amplitude. The results of this study are shown in Figure 8 (b) and (d), which show the time-dependent acceleration amplitude for both the standard ceramic tile and the metatile. The impact noise source is represented by the blue curve, while the transmitted noise is shown in red. Standard ceramic tile exhibits high transmitted noise, characterized by a large amplitude over time. In contrast, the metatile shows a significantly lower amplitude, indicating its effectiveness in attenuating noise propagation.
@@
To further analyze this behavior, the Fast Fourier Transform (FFT) of both responses is calculated and presented in Figure 8(e) and (f). The standard ceramic tile allows noise transmission across the frequency spectrum from $20$ Hz to $2000$ Hz. However, the metatile demonstrates a significant reduction in noise transmission at low frequencies, and beyond $500$ Hz, no impact noise transmission is observed. These results highlight the strong noise control and attenuation capabilities of the metatile, making it a promising solution for tiled flooring applications.
The comparison between the standard ceramic tile and our newly proposed design, referred to as the ceramic metatile, demonstrates that the inclusion of silicon joints significantly enhances the elastodynamic behavior of the tile. This enhancement is evident in the emergence of bandgaps and the improved attenuation of impact noise on tiled floors. Notably, the materials used in the ceramic metatile are the same as those in standard ceramic tiles, and the use of silicon adhesive is already common in real-world applications for joining tiles. Our study shows that, by combining these existing materials with structural engineering principles, it is possible to substantially improve impact noise and vibration comfort in tiled rooms.
In terms of cost and material effectiveness, the proposed ceramic metatile remains comparable to standard ceramic tiles. The only modification involves using sub-tiles with dimensions of $5\times 30cm^{2}$. These can be produced either by designing a new mold or by cutting existing standard tiles using a ceramic cutter. The silicon joints, already used in tile installations, can be repurposed in the same manner for constructing the metatile. Regarding long-term durability, the metatile is expected to perform similarly to standard vitrified ceramics, which are well-known for their long-lasting performance. As for installation, the process is straightforward and compatible with conventional methods. One approach involves placing the sub-tiles individually and applying the silicon joints between them to form the metatile directly on-site. Alternatively, preassembled ceramic metatiles could be manufactured and installed using magnetic backing sheets, offering easy installation and replacement. In summary, our design not only exhibits strong attenuation of elastic waves but also holds promise for further optimization to target lower frequencies, potentially below $100$ Hz, while maintaining compatibility with current building practices and materials.
Additionally, further design strategies based on the use of silicon joints or other soft materials can enhance the insulation performance of the tiled floor. In this study, we presented a unidirectional arrangement of the sub-tiles and silicon joints. However, an alternative configuration is to arrange the sub-tiles and silicon joints along both the $x$- and $y$-directions (see Appendix C), which could further improve attenuation in both directions of the tiled floor.
5 Conclusion
To conclude, we presented novel designs of ceramic metatiles that combine ceramic tiles with silicon joints. To characterize their elastic insulation behavior, we performed numerical optimizations and calculated their transmission and dispersion curves in order to optimize the dynamic response of the designs. Subsequently, we validated our numerical approach through experimental testing. For that, we fabricated homemade ceramic metatiles and employed a laboratory setup using Fast Fourier Transformation to measure their transmission responses. Furthermore, we highlighted the standard acoustic properties of standard ceramic tiles, commonly used in tiled floors, as a reference to contrast with the outcomes obtained from our ceramic metatiles. As a founding, we found broadband bandgaps at low-frequency ranges for both flexural and longitudinal waves, in contrast to standard tiles where such features are less pronounced for elastic waves. Our findings showed that for flexural waves, the ceramic metatile exhibits a significant reduction in transmission within the 5005005005001900190019001900 Hz range along the $[100]$ direction and within the 5005005005001400140014001400 Hz range along the $[110]$ direction. For longitudinal waves, the ceramic metatile displays a broad bandgap extending from $400$ Hz to $1950$ Hz along both the $[100]$ and $[110]$ directions. In addition, we observed a shift of the bandgaps toward lower frequencies as the width of the subtiles and silicon joints increases, or as the Young’s modulus of silicon decreases. This indicates that the stop bands can be tailored through geometric parameters and material properties to effectively target low-frequency ranges, such as those associated with footstep and indoor noise propagating through elastic waves in tiled floors. Finally, the homemade ceramic metatile samples confirmed the numerical results, the experimental and numerical data are in good agreement, both showing strong elastic stop bands for elastic impact noise. This validates the efficiency and reliability of our design in attenuating elastic wave propagation through tiled floors. Overall, we demonstrated the potential of our novel ceramic metatile design for practical applications in architectural acoustics, particularly in controlling unwanted vibrations and noise in building indoors.